how would i make graphs for each answer?

1. Write the equation of the line parallel to the line x + 4y = 5 with a y-intercept at ( 0,-2 ). Write your answer in standard form.
x + 4y = 5
-x -x
--------------
4y= -x +5
4y/4 -x/4 5/4
y = -1/4 + 5/4
the slope is -1/4
y = -1/4x - 2
+1/4 +1/4
------------------
1/4x + y = -2
x + 4y = -8



2.Write the equation of the line parallel to the line 3x - 4y = 7 and contains the point (8,-1). Write your answer in standard form.
3x - 4y = 7
-3x -3x
---------------
-4y = -3x + 7
-4y/4 -3x/-4 7/-4
y = 3/4x - 7/4
the slope is 3/4
y - (-1) 3/4
3x -4 =28



3. Write the equation of the line perpendicular to the line 4x - 5y = 1 and passes through the point (-8,5). Write your answer in standard form.
4X - 5Y = 1
-4X -4X
--------------
-5Y = -4 + 1
-5Y/-5 -4/-5 1/-5
y = 4/5x - 1/5
slope is -5/4
y - 5 = -5/4 ( x - 8)
y - 5 = -5/4x -10
y = -5/4x - 5
5/4x + y = -5
4 (5/4x +y =-5)
5x + 4y = -20



4. Write the equation of the line perpendicular to the line 2x + 7y = 3 and passes through the point (0,6). Write your answer in standard form.
2x + 7y = 3
-2x -2x
---------------
7y= -2/7x + 3
7y/7 -2x/7 3/7
the slope is 7/2
y = 7/2x + 6
7/2x + 7 = 6
2 (-7/2x + y =6 )
-7x + 2y = 12
7x - 2y = -12

Answer by Alan3354(31538)

how would i make graphs for each answer?
1. Write the equation of the line parallel to the line x + 4y = 5 with a y-intercept at (0,-2). Write your answer in standard form.
x + 4y = 5
-x -x
--------------
4y= -x +5
4y/4 -x/4 5/4
y = -1/4 + 5/4
the slope is -1/4
You're correct to this point. Then, to find the eqn with the slope and a point, use:
y-y1 = m*(x-x1) where (x1,y1) is the point.
----------------
y - (-2) = (-1/4)*(x-0)
y+2 = -x/4
y = (-1/4)x - 2 (slope-intercept form)
4y = -x-8
x + 4y = -8 (standard form)
---------------
---------------
y = -1/4x - 2
+1/4 +1/4
------------------
1/4x + y = -2
x + 4y = -8


2.Write the equation of the line parallel to the line 3x - 4y = 7 and contains the point (8,-1). Write your answer in standard form.
3x - 4y = 7
-3x -3x
---------------
-4y = -3x + 7
-4y/-4 = -3x/-4 + 7/-4
y = 3/4x - 7/4
the slope is 3/4
----------------
Same as the one above from this point.
------------------
y - (-1) 3/4
3x -4 =28


3. Write the equation of the line perpendicular to the line 4x - 5y = 1 and passes through the point (-8,5). Write your answer in standard form.
4X - 5Y = 1
-4X -4X
--------------
-5Y = -4 + 1
-5Y/-5 -4/-5 1/-5
y = 4/5x - 1/5
slope is -5/4
y - 5 = -5/4 ( x - 8)
y - 5 = -5/4x -10
y = -5/4x - 5
5/4x + y = -5
4 (5/4x +y =-5)
5x + 4y = -20


4. Write the equation of the line perpendicular to the line 2x + 7y = 3 and passes through the point (0,6). Write your answer in standard form.
2x + 7y = 3
-2x -2x
---------------
7y= -2/7x + 3 **** check this again
7y/7 -2x/7 3/7
the slope is 7/2
y = 7/2x + 6
7/2x + 7 = 6
2 (-7/2x + y =6 )
-7x + 2y = 12
7x - 2y = -12
------------------
If you have questions, or want me to check your work, email me via the thank you note.